Circle the quantity that maximizes total revenue (TR) for the marginal revenue (MR) function Q = 75 – 7P

A.38.25 B. 44.48 C. 49.41 D. 50.50 E. 59.30 F. 75.00
Answer #1

marginal Revenul Integrating 7 PQ = 750 Funchich 4 = 75-7P 7 P = 75-4 Both sides 2 2 7 TR = 75U- 2 TR = 7 Max TR 2) dTR ZO dy